### v1

@@ Abstract
Multistakeholder recommender systems are those that account for the impacts and preferences of multiple groups of individuals, not just the end users receiving recommendations. Due to their complexity, evaluating these systems cannot be restricted to the overall utility of a single stakeholder, as is often the case of more mainstream recommender system applications. In this article, we focus our discussion on the intricacies of the evaluation of multistakeholder recommender systems. We bring attention to the different aspects involved in the evaluation of multistakeholder recommender systems—from the range of stakeholders involved (including but not limited to producers and consumers) to the values and specific goals of each relevant stakeholder. Additionally, we discuss how to move from theoretical principles to practical implementation, providing specific use case examples. Finally, we outline open research directions for the RecSys community to explore. We aim to provide guidance to researchers and practitioners about how to think about these complex and domain-dependent issues of evaluation in the course of designing, developing, and researching applications with multistakeholder aspects.
@@ 1 Introduction
To develop a holistic view of the operation of a recommender system, it is often important to consider the impact of the system beyond just the primary users who receive recommendations – although the perspectives of such users will always have a primary role in a personalized system. Expanding the frame of evaluation to include other parties, as well as the ecosystem in which the system is deployed, leads us to a multistakeholder view of recommender system evaluation as defined in [2]:
@@
”A multistakeholder evaluation is one in which the quality of recommendations is assessed across multiple groups of stakeholders.”
@@
In this article, we provide (i) an overview of the types of recommendation stakeholders that can be considered in conducting such evaluations, (ii) a discussion of the considerations and values that enter into developing measures that capture outcomes of interest for a diversity of stakeholders, (iii) an outline of a methodology for developing and applying multistakeholder evaluation, and (iv) three examples of different multistakeholder scenarios including derivations of evaluation metrics for different stakeholder groups in these different scenarios.
@@
The variety of possible stakeholders we identified that are part of the general recommendation ecosystem is suggested in Figure 1 and defined here, using the terminology from [1, 2]:
@@
    are the traditional recommender system users to whom recommendations are delivered and to which typical forms of recommender system evaluation are oriented.
@@
    are those who are impacted by choices that recommendation consumers make, by interacting with chosen items or being impacted by the use or consumption of recommended items. For example, in a recommender system that suggests children’s books to teachers, the children who ultimately get the books (and their parents) are downstream stakeholders from the teachers who are users of the system [15, 18].
@@
    is intended to stand in for the organization creating and operating the recommendation platform itself. This group may have a variety of values, including, but not limited to, economic ones that are not necessarily shared by the consumers or providers.
@@
    are those individuals or groups who do not have direct interaction with the system that nonetheless have an interest or are impacted by its operation. For example, in a domain such as job recommendation, government agencies charged with ensuring non-discrimination in hiring practices may be considered stakeholders whose requirements are legally binding on the platform operator.
@@
The vast majority of recommender systems research focuses its evaluation only on the perspective of recommendation consumers. However, in most applications, numerous stakeholders are involved in the upstream and downstream parts of the provisioning, recommending, and consumption process. Here, we illustrate this complexity using a (hypothetical) music streaming application as an example—additional examples from other application areas are described in Section 4.
Figure 2 shows the different stakeholders involved in the process, with songwriters, artists, and label companies on the content production and provisioning side. The platform (recommender system) plays the role of mediating between upstream and downstream stakeholders. On the downstream side, consumers are the first-line stakeholders, but others may also be affected by the recommendations, e.g., owners of concert venues where recommended artists might appear.
Figure 2: Stakeholder relations for the music streaming example.
Stakeholders pursue specific goals that are driven by values (see Section 2) meant as generic concepts helping an individual to choose the best actions or behaviors [61, 63]. While values are generic concepts and may apply across a wide range of application domains, goals can be seen as intermediate-level objectives that are operationalizations of, for example, a generic human- or business-centric value. Each goal can be assessed by different measures, which may be captured using a variety of concrete measurement methods and metrics [19]. Unsurprisingly, the goals of different stakeholders may compete with each other, creating the need to balance stakeholder goals in the recommendation process. For the music streaming example, we present sample goals and measures in Table 1. In this context, conflicting goals may be that system operators want to increase the monetary benefit by favoring popular artists and songs which might negatively affect the visibility of long-tail artists who want to build an audience. [^1] We stress that all examples in this discussion are hypothetical and may or may not represent actual stakeholder configurations or goals. For additional perspectives on multi-objective recommendation in music recommendation, see [72].
@@
Stakeholder Artist / Songwriter Music Label Streaming Service Listener Concert Venue
@@
Customer loyalty
Enjoyment, Wellbeing,
Personal development
Monetary reward,
Market development,
@@ 1.2 Challenges
  • Application specificity: Recommender systems research is, in general, highly domain specific. This specificity is even more pronounced when the larger ecosystem is considered. As our examples make clear (see Section 4), different recommendation applications have different stakeholder configurations and different types of benefits of utility that stakeholders may gain. Even across recommender applications for which outputs are superficially similar (for example, music playlists), stakeholders may occupy different niches and require different analyses.
  • Access to data: Typical recommendation datasets have little to no information about non-consumer stakeholders, so it is difficult to understand what are realistic calculations of, for example, revenue distribution among item providers. Additional work will typically be required to gather the data needed to design effective multistakeholder evaluations.
@@
  • Adversarial aspects: Recommendation platforms may actively discourage providers in particular from acquiring knowledge about the platform that might enable strategic activity: for example, misrepresenting their items to gain algorithmic favor. There is no doubt that providers are sometimes incentivized to do this, as the history of search engine spam attests. It is an open research question to design evaluation metrics that can be shared with providers without enabling adversarial behavior.
@@ 2 Values
Jannach and Zanker [31] mention that, ideally, recommender systems would “create value in parallel for all involved stakeholders”. At the same time, it is unavoidable for competing goals to arise, since direct and indirect stakeholders, including the system itself, may have their own perspectives. In this case, to evaluate the value created for those involved, we argue that it is imperative to go back to a fundamental and normative question and one that is rarely asked according to Jannach and Zanker [30]: “What is a good recommendation (in a given context)?
@@
To answer this complex question, we posit that one first must look into the values each stakeholder aims for in this multistakeholder process. The concept of ‘value’ has been discussed in the literature from multiple perspectives [28, 70, 2, 9, 69, 25, 53, 69]. Perhaps the most prominent are those referring to the business side of the equation (provider-centered) or the user side (consumer-centered), i.e., the utility of the ultimate consumer. From a more human perspective, values concerning individuals directly or indirectly served by recommender systems and those with societal implications have also been discussed. However, as seen in various practical applications of multistakeholder recommendation tasks, this concept can often be open to multiple interpretations.
@@
In the context of this work, we refer to value as standards or criteria that help an individual to select and evaluate actions or behaviors [61, 63]. With that in mind, for multistakeholder recommender systems, the term value might refer to standards (or even a set of standards) a stakeholder expects or imposes on the recommendation process. The significance of values in system design has been highlighted within the field of human-computer interaction through the development of value-sensitive design processes [50, 23]. In multistakeholder recommender systems, values must be considered when evaluating the ‘goodness’ not just of a recommendation itself, but of the stakeholders that are part of the entire process within the specific contexts and domains in which the systems are deployed.
@@
In the remainder of this section, we review seminal literature that provides background on the concept of value from different perspectives and its connection to recommender systems. Along the way, we highlight the most common values to reflect on when evaluating multistakeholder recommender systems. It is worth noting that the values we mention are not meant to be an exhaustive list. Instead, they serve as a starting point to encourage reflection among researchers and practitioners to move beyond the more typical ‘producer versus consumer’ perspective and consider a myriad of factors to (simultaneously) account for when evaluating multistakeholder recommender systems.
@@ 2.1 Economic and Business-Related Values
When addressing values in the context of multistakeholder recommender system evaluation, economic and business-related values are often considered, especially for providers and system operators.
@@
De Biasio et al. [9] provide a systematic review of value-aware recommender systems, introducing value primarily as an economic concept leading to monetary reward (i.e., profit and revenue). They distinguish several aspects that inform the value of monetary reward reflective of a business and economic view, including use value (e.g., increasing revenue by providing useful recommendations), estimated value (related to attractiveness and desirability, such as having a comprehensive music catalog to create recommendations from), cost value (e.g., the economic resources required to distribute a music album to the music streaming platform), and exchange value (the change in value over time, e.g., increase in a music artist’s recognition and popularity on the platform due to effective recommendations).
@@
From this, we observe values related to user perception and customer loyalty, which are crucial from both a business and economic perspective. These values often relate to “the concepts of quality and personalization, experience and trust, features, and benefits” [9]. For example, in the music industry, a platform that provides highly personalized playlists based on users’ listening history can significantly enhance user satisfaction. This personalization not only helps users discover new music that aligns with their preferences but also fosters a sense of trust and loyalty towards the platform. Users are more likely to stay subscribed and recommend the service to others if they consistently experience high-quality, relevant recommendations.
@@
The authors in [10] highlight that recommender systems typically serve an organization’s economic values. Besides profit and revenue (i.e., monetary rewards), this might be related to growth and market development. For example, music streaming platforms aim to generate profit and attract new users by offering social features like joint playlist creation, which benefit users when their peers are also on the platform. Furthermore, the authors characterize economic recommender systems as systems that exploit “price and profit information and related concepts from marketing and economics to directly optimize an organization’s profitability.Jannach and Adomavicius [28] identify strategic perspectives for both consumers and providers. For consumers, personal utility includes happiness, satisfaction, knowledge, and entertainment. For providers, organizational utility encompasses profit, revenue and growth. In addition, other values, such as changing user behavior to create demand, might be relevant. For example, a music streaming platform might recommend emerging artists or newly released tracks to users, encouraging them to explore and adopt new music preferences, thereby creating demand for content that the platform can better monetize.
@@
Jannach and Zanker [31] examine the theory of business models in e-commerce recommender systems and identify the following value-driving aspects: efficiency (e.g., the exposure of music artists in recommendation lists or the number of clicks on recommended music tracks), complementarities (e.g., creating value through synergies by combining different item types like recommending merchandise articles along with track recommendations of a specific music artist), lock-in and churn prevention (e.g., retaining subscribed users by providing meaningful recommendations), and novelty and product planning (e.g., finding new fans through recommendations to users who might like an artist’s music or getting inspired to create new music album).
@@
In addition to immediate financial outcomes, recommender systems can enhance a platform’s brand equity by creating positive user experiences that increase user satisfaction and loyalty, both critical components of brand equity. Jannach and Jugovac [29] emphasize that well-crafted recommendations enable platforms to differentiate themselves in competitive markets, thereby strengthening their reputation. Similarly, Maslowska et al. [45] suggest that when recommendations align with users’ personal goals, they not only encourage engagement but also create a positive spillover effect on the platform’s brand, enhancing overall trust and loyalty. Recommender systems also open up cross-selling and up-selling opportunities in ways that reinforce brand value, such as suggesting premium products or exclusive experiences that align with the brand’s identity (e.g., premium subscriptions or concert tickets). Finally, platforms can also use user interaction data from recommender systems to develop personalized marketing strategies while maintaining high privacy standards, a practice that reinforces consumer trust and encourages long-term engagement [45].

### v2

@@ Abstract
Multistakeholder recommender systems are those that account for the impacts and preferences of multiple groups of individuals, not just the end users receiving recommendations. Due to their complexity, these systems cannot be evaluated strictly by the overall utility of a single stakeholder, as is often the case of more mainstream recommender system applications. In this article, we focus our discussion on the challenges of multistakeholder evaluation of recommender systems. We bring attention to the different aspects involved—from the range of stakeholders involved (including but not limited to providers and consumers) to the values and specific goals of each relevant stakeholder. We discuss how to move from theoretical principles to practical implementation, providing specific use case examples. Finally, we outline open research directions for the RecSys community to explore. We aim to provide guidance to researchers and practitioners about incorporating these complex and domain-dependent issues of evaluation in the course of designing, developing, and researching applications with multistakeholder aspects.
@@ 1 Introduction
Understanding the performance of any machine learning system requires evaluation. Research in recommender systems has focused almost exclusively on evaluating the experience of the end user receiving recommendations, even though recent research highlights the crucial role that recommender systems play in larger ecosystems of commerce and media distribution [3, 11]. To encompass the larger context of a complex recommender ecosystem requires taking a broader view of evaluation (although end users will always be important). This brings us to the topic of multistakeholder evaluation as defined in [3]:
@@
“A multistakeholder evaluation is one in which the quality of recommendations is assessed across multiple groups of stakeholders.”
@@
The aim of this article is to provide an introduction to the topic of multistakeholder evaluation of recommender systems. We provide (i) an overview of the types of recommendation stakeholders that may be considered in conducting such evaluations, (ii) a discussion of the considerations and values that enter into developing measures that capture outcomes of interest for a diversity of stakeholders, (iii) an outline of a methodology for developing and applying multistakeholder evaluation, and (iv) three examples of different multistakeholder scenarios including derivations of evaluation metrics for different stakeholder groups in these scenarios. While it goes without saying that a system designed with multiple stakeholders in mind will need to be evaluated relative to the perspectives of such stakeholders, multistakeholder evaluation does not necessarily imply that the recommender system in question accounts for multiple stakeholders in its design or optimization objectives. One may wish to audit the performance of the system relative to different stakeholders interests without necessarily feeding all such measures back to the algorithm.
This work derives from a Dagstuhl seminar titled Evaluation Perspectives of Recommender Systems: Driving Research and Education [119]. [^1] https://www.dagstuhl.de/en/seminars/seminar-calendar/seminar-details/24211, May 2024 This article reflects the perspectives of a sub-group of the researchers in attendance focused on the multistakeholder evaluation of recommender systems [14]. These researchers individually contribute knowledge in multistakeholder methodologies, their evaluation, as well as application to several domains (music, HR, and education). In addition, the cumulative experience of these researchers points to gaps in the multistakeholder evaluation for recommender systems. Notably, we argue that the research to date does not have a clear method for developing evaluations that reflect stakeholder impact or the wider range of possible impacts. In this article, we make a case for revisiting value-sensitive design (c.f., [38]) and indicate necessary steps for developing evaluation protocols, in addition to settling on metrics, and handling trade-offs in stakeholder preferences. Finally, we argue that multistakeholder evaluation necessarily reflects (often implicit) values (c.f., [105]), and that impactful evaluation requires the identification not only of a wider range of stakeholders, but also the specific characteristics of a task such as: application specificity, context specificity, and institutional sensitivity.
Evaluating a methodology aimed at the development of software applications usually goes through the definition of three main components [58]: (i) a list of feature requirements, (ii) a method of scoring the features in the methodologies targeted, and (iii) a set of guidelines for applying the evaluation framework. Among these components, a first-class role is played by the identification of relevant features to assess the evaluation. A key element for the evaluation is the identification of similarities and differences with other methodologies as well as to set a common reference. This latter can always be used to set a comparative evaluation of different methodologies as well as of different choices within the same methodology. In the multistakeholder scenario we target, a common reference baseline can always be set as the ”classical” consumer-only one.
In this work, we provide the reader with all the elements to evaluate the adopted methodology with reference to the components mentioned before.
  • Given the targeted stakeholders (a classification is provided in Section 1.1), for each of them the system designer defines a set of features that follow clear values and goals as highlighted in Section 2.
  • The selected features, referring to different stakeholders, can be measured and scored, as discussed in Section 3.3, and eventually aggregated (see Section 3.4) to provide an overall holistic evaluation of the selected features/values.
  • A set of practical guidelines are then provided in Section 3.5 for conducting multistakeholder evaluation.
In order to support the application of the multistakeholder evaluation, in Section 4 we show how to apply the above mentioned guidelines with the introduction of some example scenarios and metrics. Finally, in Section 5 we describe potential challenges in multistakeholder recommendation.
1.1 Stakeholder types
@@
The variety of stakeholders forming a recommendation ecosystem is suggested in Figure 1 and defined here, extending the terminology from [2, 3]:
@@
    are the traditional recommender system end users to whom recommendations are delivered and to which typical forms of recommender system evaluation are oriented.
@@
    are those who are impacted by choices that recommendation consumers make, by interacting with chosen items or being impacted by the use or consumption of recommended items. For example, in a recommender system that suggests children’s books to teachers, the children who ultimately get the books (and their parents) are downstream stakeholders from the teachers who are the consumers, receiving recommendations from the system [29, 32].
@@
    is intended to stand in for the organization creating and operating the recommendation platform itself. A recommender system is implemented and maintained for a reason, in support of an organization’s goals. An e-commerce site may implement a recommender system to achieve goals around user satisfaction and retention, quite related to the goals of consumers, but they may also have economic goals (increasing average purchase basket size, for example) that are not necessarily shared by the consumers or providers.
@@
    are those individuals or groups who do not have direct interaction with the system that nonetheless have an interest in, or are impacted by, its operation. For example, in a domain such as job recommendation, government agencies charged with ensuring non-discrimination in hiring practices may be considered stakeholders whose requirements are legally binding on the platform operator.
@@
As we have noted, the vast majority of recommender systems research focuses its evaluation only on the perspective of recommendation consumers. However, in most applications, numerous stakeholders are involved in the upstream and downstream parts of the provisioning, recommending, and consumption process. Here, we illustrate this complexity using a (hypothetical) music streaming application as an example—additional examples from other application areas are described in Section 4.
Figure 2 shows the different stakeholders involved in the system, with songwriters, artists, and label companies on the content production and provisioning side. The platform (recommender system) plays the role of mediating between upstream and downstream stakeholders. On the downstream side, consumers are the first-line stakeholders, but others may also be affected by the recommendations, e.g., owners of concert venues where recommended artists might appear.
Figure 2: Stakeholder configuration for the music streaming example.
Stakeholders pursue specific goals that are driven by values (see Section 2) meant as generic concepts helping an individual to choose the best actions or behaviors [96, 98]. While values are generic concepts and may apply across a wide range of application domains, goals can be seen as intermediate-level objectives that are operationalizations of, for example, a generic human- or business-centric value. Each goal can be assessed by different measures, which may be captured using a variety of concrete measurement methods and metrics [33]. For the music streaming example, we present sample goals and measures in Table 1; these are described in greater detail in rest of the article.
Unsurprisingly, the goals of different stakeholders may compete with each other. A system that is seeking to satisfy competing stakeholders will need to balance stakeholder goals in the recommendation process. In this example, conflicting goals may be that system operators want to increase the monetary benefit by favoring popular artists and songs, which might negatively affect the visibility of long-tail artists who want to build an audience. [^2] We stress that all examples in this discussion are hypothetical and may or may not represent actual stakeholder configurations or goals. For additional perspectives on multi-objective recommendation in music recommendation, see [112].
@@
Stakeholder
Artist /
Songwriter
Music
Label
Streaming
Service
Listener
Concert
Venue
@@
Monetary
reward,
Customer loyalty
Enjoyment,
Wellbeing,
Personal
development
@@
Market
development,
@@ 1.3 Challenges
  • Application specificity: Recommender systems research is, in general, highly domain specific. This specificity is even more pronounced when the larger ecosystem is considered. As our examples make clear (see Section 4), different recommendation applications have different stakeholder configurations and different types of benefits that stakeholders may gain. Even across recommender applications for which outputs are superficially similar (for example, music playlists), stakeholders may occupy different niches and require different analyses.
  • Access to data: Typical recommendation datasets have little to no information about non-consumer stakeholders, so it is difficult to understand what are realistic calculations of, for example, revenue distribution among item providers. Additional investigation will typically be required to gather the data needed to design effective multistakeholder evaluations.
@@
  • Adversarial aspects: Recommendation platforms may actively discourage providers in particular from acquiring knowledge about the platform that might enable strategic activity: for example, misrepresenting their items to gain algorithmic favor. There is no doubt that providers are sometimes incentivized to do this, as the history of search engine spam attests. Although adversarial considerations do not limit a platform’s ability to conduct internal multistakeholder evaluation, it is an open research question how to design evaluation metrics and disclosure protocols such that outcomes can be shared with providers without enabling adversarial behavior.
1.4 Related Work
Multistakeholder recommendation has long been a subject of study in the field although the term itself was not coined until 2017 [2]. The (pre)history of multistakeholder recommendation is detailed by Abdollahpouri et al. [3]. In these earlier works, researchers defined evaluation approaches deemed appropriate to their particular task, measuring the outcomes important to the aspects of recommendation under consideration. For example, early work by Chen et al. [18] combined profitability with standard measures of recall and precision in evaluating a product recommender. However, because multistakeholder recommendation was not conceptualized as a research area, connections between these research areas and a general theory of multistakeholder evaluation in recommender systems did not materialize.
Abdollahpouri et al. [3] call out multistakeholder evaluation specifically as a research topic in their survey. They specifically formulate a number of possible provider metrics. However, unlike our approach here, this work is not grounded in value-centric consideration of stakeholder perspectives. They call for practitioners to formulate metrics from the point of view of different stakeholders but do not consider making use of stakeholder involvement in the process. Similarly, in the area of fairness-aware recommendation, a subtype of multistakeholder recommendation, Ekstrand et al. [31] survey a wide variety of fairness metrics that have been used in the literature. However, these metrics are generally not the product of stakeholder consultation. Some steps in this direction are taken in [101] and [102] where stakeholder consultation in the form of qualitative interviews is used to derive fairness concerns and possible operationalizations. Multistakeholder aspects of machine learning fairness in industry practice more generally are explored in [103]. This work finds that practitioners have little leeway to conduct broad stakeholder consultation in most circumstances, even when they believe this may be valuable. Practitioners report that they have little to go on when trying to understand and operationalize objectives representing different stakeholder interests.
In general, the Dagstuhl seminar participants found that there was insufficient guidance available on how to design and conduct multistakeholder evaluation of recommender systems. In particular, we believe that evaluation must be rooted in understanding what stakeholders value in a system and that this step of analysis is treated implicitly or not at all in most prior work.
@@ 2 Values
Jannach and Zanker [55] argue that, in the ideal case, recommender systems would “create value in parallel for all involved stakeholders”. At the same time, it is unavoidable for competing goals to arise, since direct and indirect stakeholders, including the system itself, may have differing perspectives. In this case, to evaluate the value created for those involved, we argue that it is imperative to go back to a fundamental and normative question and one that is rarely asked according to Jannach and Zanker [54]: “What is a good recommendation (in a given context)?
@@
To answer this complex question, we posit that one first must look into the values each stakeholder aims for in the recommendation process. The concept of value has been discussed in the recommender systems literature from multiple perspectives [50, 110, 3, 21, 109, 43, 87]. Perhaps the most prominent are those referring to the business side of the equation (provider-or system-centered) or the user side (consumer-centered), i.e., the utility of the ultimate consumer. From a more human perspective, values concerning individuals directly or indirectly served by recommender systems [108, 109, 62, 28], and those with societal implications [34, 80, 81, 45, 109] have also been discussed, such as the level of social utility that it achieves, the social rights that it strengthens or harms, and the existing social norms that it alters or reinforces. The concept of values in domain-specific implementations of recommender systems has also been investigated, e.g., news recommender systems [6] or recommender systems for humanities and historical research [5]. Although these works analyzed a specific domain for recommender systems, we find that many of the identified values overlap and that both more business-oriented and more human-oriented values are discussed.
@@
In the context of this work, we refer to value as standards or criteria that help an individual to select and evaluate actions or behaviors [96, 98]. With that in mind, for multistakeholder recommender systems, the term value might refer to standards (or even a set of standards) a stakeholder expects or imposes on the recommendation process. The significance of values in system design has been highlighted within the field of human-computer interaction through the development of value-sensitive design processes. We point the interested reader who seeks a deeper understanding on value-sensitive design for information systems to the literature in the field [82, 40, 39, 10].
In multistakeholder recommender systems, values must be considered when evaluating the ‘goodness’ not just of a recommendation in isolation, but of the stakeholders that are part of the entire process within the specific contexts and domains in which the systems are deployed. Thus, these systems can be conceptualized as complex socio-technical systems in which algorithmic processes are embedded within broader structures of negotiation, deliberation, and governance. Particularly in public contexts, values are not merely elicited from stakeholders but are shaped through deliberative processes that mediate trade-offs and articulate shared goals. This perspective highlights the importance of viewing value alignment not just as a technical optimization problem, but as a process of participatory design. Insights from public policy and political science, such as theories of deliberative democracy, stakeholder governance, and collaborative planning, can therefore enrich the value elicitation processes of multistakeholder recommender systems by grounding them in frameworks that acknowledge power asymmetries, legitimacy, and fairness. We therefore, also point the interested reader to the rich body of literature in this realm [e.g., 25, 48, 107].
Value-associated considerations in technology have in recent years also become subject of political discourse and policy making. The European Union, for instance, has issued legal regulations that aim at safeguarding against adverse effects of digital services in general and, specifically, AI-based systems. The Digital Services Act  [^4] https://eur-lex.europa.eu/legal-content/EN/TXT/?uri=CELEX3A32022R2065 (DSA) addresses a wide range of online platforms, generally setting the goal of “ensuring a safe, predictable and trustworthy online environment”. The DSA repeatedly spotlights recommender systems, requesting, inter alia, that “recipients of their service are appropriately informed about how recommender systems impact the way information is displayed, and can influence how information is presented to them”. Providers should also clearly present the parameters, at least the most important criteria, to let users understand how the information is prioritized for them. The impact of the DSA on multistakeholder recommender systems has already become subject of scientific discourse (see, e.g., [89]).
Specific regulations were recently also adopted for AI-based systems  [^5] https://eur-lex.europa.eu/eli/reg/2024/1689/oj/eng (the EU AI Act). The approach taken there is risk-based, classifying applications into several risk classes. While many recommender systems (as one relevant class of AI-based systems) can be assumed to involve only acceptable risk for stakeholders, e.g., in music streaming, other domains addressing, e.g., the job market or medical applications may indeed create substantial risks for some stakholder groups [66]. Identifying such risks should therefore become integral part of the value elicitation process for stakeholder groups that may be critically affected. Both the DSA and the AI Act are relevant for different stakeholder groups, but do not specifically address multistakeholder issues.
@@
In the rest of this section, we review seminal literature that provides background on the concept of value from different perspectives and its connection to recommender systems. Along the way, we highlight the most common values to consider when evaluating multistakeholder recommender systems. It is worth noting that the values we mention are not meant to be an exhaustive list. Instead, they serve as a starting point to encourage reflection among researchers and practitioners to move beyond the more typical ‘producer versus consumer’ perspective and consider a myriad of factors to (simultaneously) account for when evaluating multistakeholder recommender systems.
@@ 2.1 Economic and Business-Related Values
When addressing values in the context of multistakeholder recommender system evaluation, economic and business-related values are often considered, especially for providers and system operators. De Biasio et al. [21] provide a systematic review of value-aware recommender systems, introducing value primarily as an economic concept leading to monetary reward (i.e., profit and revenue). They distinguish several aspects that inform the value of monetary reward reflective of a business and economic view, including use value (e.g., increasing revenue by providing useful recommendations), estimated value (related to attractiveness and desirability, such as having a comprehensive music catalog to create recommendations from), cost value (e.g., the economic resources required to distribute a music album to the music streaming platform), and exchange value (the change in value over time, e.g., increase in a music artist’s recognition and popularity on the platform due to effective recommendations).
@@
From this perspective, we also observe values related to user perception and customer loyalty, which are crucial from both a business and economic perspective. These values often relate to “the concepts of quality and personalization, experience and trust, features, and benefits” [21]. For example, in the music industry, a platform that provides highly personalized playlists based on users’ listening history can significantly enhance user satisfaction. This personalization not only helps users discover new music that aligns with their preferences but also fosters a sense of trust and loyalty towards the platform. Users are more likely to stay subscribed and recommend the service to others if they consistently experience high-quality, relevant recommendations.
@@
The authors in [22] highlight that recommender systems typically serve an organization’s economic values. Besides profit and revenue (i.e., monetary rewards), this might be related to growth and market development. For example, music streaming platforms aim to generate profit and attract new users by offering social features like joint playlist creation, which benefit users when their peers are also on the platform. Furthermore, the authors characterize economic recommender systems as systems that exploit “price and profit information and related concepts from marketing and economics to directly optimize an organization’s profitability.Jannach and Adomavicius [50] identify strategic perspectives for both consumers and providers. For consumers, personal utility includes happiness, satisfaction, knowledge, and entertainment. For providers, organizational utility encompasses profit, revenue and growth. In addition, other values, such as changing user behavior to create demand, might be relevant. For example, a music streaming platform might recommend emerging artists or newly released tracks to users, encouraging them to explore and adopt new music preferences, thereby creating demand for content that the platform can monetize.
@@
Jannach and Zanker [55] examine the theory of business models in e-commerce recommender systems and identify the following value-driving aspects: efficiency (e.g., the exposure of music artists in recommendation lists or the number of clicks on recommended music tracks), complementarities (e.g., creating value through synergies by combining different item types like recommending merchandise articles along with track recommendations of a specific music artist), lock-in and churn prevention (e.g., retaining subscribed users by providing helpful recommendations), and novelty and product planning (e.g., finding new fans through recommendations to users who might like an artist’s music or getting inspired to create new music album).
In addition to immediate financial outcomes, recommender systems can enhance a platform’s brand equity, which refers to the value added to a product or service by its brand as perceived by consumers. Brand equity is built through positive user experiences that increase user satisfaction and loyalty, both critical components of brand equity [1, 114]. Jannach and Jugovac [53] emphasize that well-crafted recommendations enable platforms to differentiate themselves in competitive markets, thereby strengthening their reputation.
@@
Similarly, Maslowska et al. [74] suggest that when recommendations align with users’ personal goals, they not only encourage engagement but also create a positive spillover effect on the platform’s brand, enhancing overall trust and loyalty. Recommender systems also open up cross-selling and up-selling opportunities in ways that reinforce brand value, such as suggesting premium products or exclusive experiences that align with the brand’s identity (e.g., premium subscriptions or concert tickets). These strategies, which refer to offering related products or higher-end options to users, help increase sales volume and raise average revenue per user [53, 67, 116]. Finally, platforms can also use user interaction data from recommender systems to develop personalized marketing strategies while maintaining high privacy standards, a practice that reinforces consumer trust and encourages long-term engagement [74].